• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Beck

    Log entry time 16:11:28 on July21,2006

    Entry number 174728

    keyword=Day shift

    8:00 - Still no beam to Hall A.
    9:00-10:00 - Controlled access:Target people were down in the hall. They fixed the temperature control of the hydrogen cell.
    10:10 - Asked for beam permit.
    ~10:15 - there is a beam permit (but still no beam)
    11:15 - asked for controlled access for changing the cables to the BCMs.
    11:30 - ask for a beam permit
    11:40 - still in controlled access, asked for another 15 min.
    12:10 - Back to beam permit after a problem with the BCMs was fixed (see hallalog)

    1:10 - MCC tries to put beam into the hall...
    4:00 - No beam until now.